OpenStudy (anonymous):

Evaluate cot870 degrees.

OpenStudy (anonymous):

the 870 degrees is too large value so you have to make it less than or equal to 360 degrees first, start by subtracting 360 from 870 unless you get a value less than or equal to 360

OpenStudy (anonymous):

\[\cot 870=\cot \left( 2*360+150 \right)=\cot 150=\cot \left( 180-30 \right)=-\cot 30\] now you can solve.

OpenStudy (anonymous):

\[-\sqrt{3}\]